OpenStudy (anonymous):

whats an equation for the line having (1,-6)and(-4,4)?

OpenStudy (anonymous):

find the slope (4- -6)/(-4-1) = 10/-5 =-2 y=-2x+b substitute a point in 4=-2(-4)+b b=-4 y=-2x-4

OpenStudy (angela210793):

(x-x1)/(x2-x1)=(y-y1)/(y2-y1) all u have to do now is replace x1=1 x2=-4 y1=-6 and y2=4...and u'll get an eq with x and y only